OpenStudy (anonymous):

Find the vertex of the parabola whose equation is y = x2 - 2x - 35

OpenStudy (anonymous):

point is(1,-36)

OpenStudy (anonymous):

if \(y=ax^2+bx+c\) the the \(x\) value of the vertex is \(-\frac{b}{2a}\) and the \(y\) value is what you get when you replace \(x\) by that number

OpenStudy (anonymous):

x=-b/2a= -(-2)/2*1=1 y=1^2-1*2-35=1-2-35=-35 so, vertex= (1,-36)

OpenStudy (anonymous):

in your case it is \(x=-\frac{-2}{2\times 1}=1\)

OpenStudy (anonymous):

try it this way

OpenStudy (anonymous):

\[y+36=(x-1)^{2}\]